OpenStudy (anonymous):

a. Out of 48 job applicants, the manager chose 22 to interview. Did more than half of the applicants get an interview? b. What percent of the applicants got an interview? Round up to the nearest whole percent. a.no; 45% b.yes; 48% c.yes; 46% d.no; 46%

OpenStudy (anonymous):

22/48 = 11/24 = ?

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

It's 46% (if rounded), but is that more or less than half? ;)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

"Did more than half of the applicants get an interview?"

OpenStudy (anonymous):

yes i think

OpenStudy (anonymous):

No, it's less. 50% of something is half.

OpenStudy (anonymous):

because 46 rounded is 50

OpenStudy (anonymous):

Hmm, not really. It would depend on the precision/significant figures; to 1 significant figure it is 50. But 11/24 is about 45.83% which could feasibly be rounded to 46%; 2 significant figures are specified in the question so it's what we aim for.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

So the answer is no, more than half the applicants did not receive an interview, only 46% (or so) did.

OpenStudy (anonymous):

thanks